PLEASE help me for this :(
--
In 3 years, Alex will be 3 times as old as his sister Precy.A year ago, Alex was 7 times as old as Precy. How old are they now?
Pls. reply immediately. THAnks :))

Answer by scott8148(6628)

"In 3 years, Alex will be 3 times as old as his sister Precy" ___ a + 3 = 3(p + 3) ___ a + 3 = 3p + 9 ___ a = 3p + 6

"A year ago, Alex was 7 times as old as Precy" ___ a - 1 = 7(p - 1) ___ a - 1 = 7p - 7 ___ a = 7p - 6

substituting ___ 3p + 6 = 7p - 6 ___ 12 = 4p ___ 3 = p

substituting ___ a = 3(3) + 6 ___ a = 15